OpenStudy (anonymous):

If T -1 (x, y) (½x, y - 5), then T(x, y)

OpenStudy (amoodarya):

your question is not clear !

OpenStudy (amoodarya):

do you mean ?\[T^{-1}(x,y)=(\frac{1}{2}x,y-6) \\tT =?\]

OpenStudy (amoodarya):

\[\frac{1}{2}x=X \rightarrow x=2X\\y-5=Y \rightarrow y=Y+5 \\so\\T(x,y)=(2X,y+5)\]

OpenStudy (anonymous):

Thank youuuuu

OpenStudy (anonymous):

what about this one

OpenStudy (amoodarya):

(x,y)=(x+3,y-2) because t(0,2)=(0+3,2-2) so find T-1

OpenStudy (amoodarya):

can you do like ,previous problem ?